# Long-Horizon Planning for Multi-Agent Robots in Partially Observable Environments

### v1

@@ Abstract
The ability of Language Models (LMs) to understand natural language makes them a powerful tool for parsing human instructions into task plans for autonomous robots. Unlike traditional planning methods that rely on domain-specific knowledge and handcrafted rules, LMs generalize from diverse data and adapt to various tasks with minimal tuning, acting as a compressed knowledge base. However, LMs in their standard form face challenges with long-horizon tasks, particularly in partially observable multi-agent settings. We propose an LM-based Long-Horizon Planner for Multi-Agent Robotics (LLaMAR), a cognitive architecture for planning that achieves state-of-the-art results in long-horizon tasks within partially observable environments. LLaMAR employs a plan-act-correct-verify framework, allowing self-correction from action execution feedback without relying on oracles or simulators. Additionally, we present MAP-THOR, a comprehensive test suite encompassing household tasks of varying complexity within the AI2-THOR environment. Experiments show that LLaMAR achieves a 30% higher success rate compared to other state-of-the-art LM-based multi-agent planners.
@@ 3 Background
Problem Setting: We consider a setting where multiple robots perform a series of everyday tasks in a home-like environment, such as cleaning a room or putting groceries in the fridge. These tasks typically require long-horizon planning, involving around 100 low-level actions to reach the goal. Our objective is to compute plans for a team of robots to execute high-level language instructions, $I$. We formalize these tasks as partially observable Markov decision processes (POMDP) (Puterman, 1994; Kaelbling et al., 1998), denoted as $\langle N,\mathcal{I},\mathcal{S},\{\mathcal{O}_{i}\},\{\mathcal{A}_{i}\},\mathcal{P},\mathcal{G},T\rangle$. $N$ is the number of agents and $\mathcal{I}$ is the high-level language instruction set. Here, $s\in\mathcal{S}$ represents the joint state of all agents, and $o\in\mathcal{O}$ denotes the observation set for all agents. Particularly, $o_{i}\in\mathcal{O}_{i}$ is the observation set of agent $i$, that captures incomplete environment state information. $a\in\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\cdots\mathcal{A}_{N}$ represents the joint action space. The joint action space comprises of different categories of high-level actions $\mathcal{A}=\mathcal{A}_{NAV}\cup\mathcal{A}_{INT}\cup\mathcal{A}_{EXP}$, where $\mathcal{A}_{NAV}$ is the joint navigation action set, $\mathcal{A}_{INT}$ is the joint interaction actions which allow the agents to interact with objects, and $\mathcal{A}_{EXP}$ are the joint exploration actions which allow the agents to explore the environment. Examples of the high-level actions include PickUp(object) $\in\mathcal{A}_{INT}$ and NavigateTo(location) $\in\mathcal{A}_{NAV}$. Each high-level action is associated with a low-level primitive action (pre-trained RL, behavior cloned, or heuristic-based policy). These actions are executed synchronously by all agents at every high-level decision step. $\mathcal{P}(s^{\prime}|s,a)$ is the joint transition probability function that defines the probability of arriving at $s^{\prime}\in\mathcal{S}$ after taking joint action $a\in\mathcal{A}$ in $s\in\mathcal{S}$. $\mathcal{G}=\{g_{1},\cdots,g_{k}\}$ defines the subtasks that the agents need to perform to accomplish the language instruction task. $T$ is the length of the planning horizon.
@@
Environment: To simulate open-ended, long-horizon tasks that resemble everyday activities, we use the AI2Thor simulator (Kolve et al., 2017), which supports a diverse set of interactions and photorealistic rendering. Since our approach does not require any parametric training, it can potentially translate to other similar embodied environments like VirtualHome (Puig et al., 2018), Habitat (Savva et al., 2019; Szot et al., 2021; Puig et al., 2023), and ThreeDWorld (Gan et al., 2021), possibly extending beyond household domains. We instantiate the problem with $N$ agents cooperating to accomplish a long-horizon rearrangement task (Batra et al., 2020) in an indoor environment. The agents do not know the objects present in the environment a priori and are encouraged to explore the environment to gather more information to complete the task. We list the complete set of actions in the Appendix B. Unlike previous solutions in similar settings (Zhang et al., 2024; Kannan et al., 2023; Wang et al., 2024), we do not rely on an oracle/simulator to verify subtask completion. In prior works, a predefined conditional satisfiability check is used as subtask completion feedback.
@@ 4 Approach
Admissible Action parsing with Semantic Translation When LMs generate action plans, natural language outputs often fail to translate to executable high-level actions. This happens when the output does not match the predefined format or refers to unrecognized contextually similar objects. We use a cosine similarity method from (Huang et al., 2022), fine-tuning a pre-trained sentence-BERT (Reimers and Gurevych, 2019) to transform the free-form text into admissible high-level actions. Hyperparameters and additional details of the sentence transformer fine-tuning are provided in Appendix G.
@@ 5 Experiments
MAP-THOR: To evaluate the performance of LLaMAR and benchmark other baseline methods, we create a benchmark dataset of tasks which we call MAP-THOR (Multi-Agent Planning tasks in AI2-THOR). While Smart-LLM (Kannan et al., 2023) introduces a dataset of 36 tasks within AI2-Thor (Kolve et al., 2017) classified by complexity, their tasks are limited to single floor plans. This limitation hinders testing the robustness of planners across different room layouts. Additionally, some tasks in their dataset cannot be performed by multiple agents, regardless of task division, such as Pick up the pillow, Open the laptop to turn it on, and Turn off the lamp.
@@
  • Implicit target location, item type and quantity: Item types and their quantities along with the target location are implicitly defined. For example, Clear the floor by placing the items at their appropriate positions. The agent is expected to place items like pens, books, and laptops on the study table, and litter in the trash can.
@@
  • Transport Rate (TR): The fraction of subtasks completed within an episode, which provides a finer granularity of task completion.
@@
  • Balance (B): The ratio between the minimum and maximum number of successful high-level actions executed by any agent that contributed towards making progress in a subtask necessary for the completion of the language instruction task. We only check for a subset of high-level actions that must be executed for accomplishing critical subtasks that leads to the successful completion of the language instruction task. If each agent $i$ out of $n$ agents completes $s_{i}$ successful tasks, the balance is defined as: $B:=\frac{\min\ \{s_{1},\cdots,s_{n}\}}{\max\{s_{1},\cdots,s_{n}\}+\epsilon}$. This measures how evenly the work is distributed among agents. A balance of zero indicates at least one agent performed no successful high-level actions, while a balance of one indicates all agents performed the same number of successful high-level actions. Here $\epsilon=1e-4$ is a small number to avoid division by zero.
  • Average steps (L): The number of high-level actions taken by the team to complete the task, capped at $T=30$ in our experiments. If the task is not completed within $T$ steps, the episode is deemed a failure.
For all the metrics, we report the means along with the $95\%$ confidence interval across all the tasks. Since SR is a binomial metric, we report the Clopper-Pearson Interval as the confidence interval.
Baselines For a fair comparison with our method, we make modifications to the baselines to make them work in partially observable settings with limited reliance on the simulator. More details about implementations can be found in Appendix E.
@@
  • ReAct (Yao et al., 2023): We use a ReAct-style prompting to let the LMs reason after suggesting high-level actions and possibly suggest ways to correct for any failures.
  • SmartLLM (Kannan et al., 2023): We modify the official codebase to only include information from the local observations of the agents instead of assuming full observability.
  • CoELA (Zhang et al., 2024): We modify the list of available high-level actions to include all possible valid combinations of actions with interactable objects in the agent’s local observation. As the scene becomes more cluttered, this list and the prompt becomes combinatorially longer. In the original implementation, the list of available actions is filtered based on the feasibility of the actions as suggested by a conditional checker.
It should be noted that Act, Chain-of-Thought, ReAct and SmartLLM are all CMAS frameworks where CoELA follows the DMAS framework.
@@ 6 Results and Discussion
Choice of the underlying LM: To understand the impact of the underlying LM’s quality on decision-making, we initially experimented with different LMs. Specifically, we utilize both the language-only and vision-language models of GPT-4 (OpenAI et al., 2023), IDEFICS-2 (Laurençon et al., 2023, 2024), LLaVA (Liu et al., 2023b, a), and CoGVLM (Hong et al., 2023a). Among these, GPT-4, when used solely with text inputs, exhibits the poorest performance. This is attributed to the agents’ inability to reason about visual observations, which is particularly detrimental for the Corrector module. Substituting GPT-4V with other vision-language models results in a decline in performance (refer Table 2) and hence we use GPT-4V as the underlying VLM while comparing to the baselines.
Underlying Success Transport Coverage Balance Steps
LM Rate Rate
GPT-4 0.51 0.85 0.95 0.83 25.80
(0.36, 0.66) (0.80, 0.91) (0.91, 0.98) (0.78, 0.86) (23.72, 27.88)
LLaVA 0.54 0.84 0.91 0.75 26.21
(0.41, 0.65) (0.71, 0.90) (0.87, 0.98) (0.64, 0.83) (21.56, 28.97)
IDEFICS-2 0.57 0.86 0.94 0.78 25.27
(0.43, 0.67) (0.74, 0.91) (0.89, 0.98) (0.65,0.84) (20.14, 28.37)
CogVLM 0.61 0.89 0.95 0.80 23.21
(0.47, 0.68) (0.73, 0.95) (0.89, 0.99) (0.73, 0.86) (20.57, 26.82)
GPT-4V 0.66 0.91 0.97 0.82 21.87
(0.50, 0.76) (0.81, 0.96) (0.93,0.99) (0.75, 0.87) (18.76, 26.43)
Table 2: Metrics by varying the underlying LLM/VLM in LLaMAR for the 2-agent scenario.
@@
Baseline Comparisons: Table 3 compares our method, LLaMAR, with other baselines in a 2-agent scenario using GPT-4V as the underlying VLM. Act and ReAct show similar performance, with Act struggling due to its lack of strategic planning or correction and ReAct performing slightly better by dynamically adjusting actions based on reasoning on immediate feedback. CoT’s performance declines with longer planning horizons due to its inability to maintain coherence over extended planning sequences, consistent with findings in (Stechly et al., 2024), showing its effectiveness only with highly specific prompts. SmartLLM, operating in a plan-and-execute paradigm, generates impractical plans with issues like infinite loops and failure to handle low-level action failures, leading to lower success rates and poor transport metrics. It also tends to hallucinate objects. CoELA, using a decentralized multi-agent system (DMAS), performs poorly due to large input prompts and struggles to select the correct action from numerous choices. Its decentralized decision-making is less efficient than the centralized multi-agent system (CMAS) used by LLaMAR. Previous research (Chen et al., 2023) confirms CMAS frameworks are more effective than DMAS frameworks. Overall, our method, LLaMAR, benefits from its modular cognitive architecture, which integrates planning, acting, correcting, and verifying through distinct LLM roles, resulting in superior performance across various evaluation metrics. By avoiding reliance on privileged information and incorporating a robust exploration strategy allowing it to scout for objects not initially visible, LLaMAR ensures higher success rates and balanced task execution among agents.
Algorithm Success Rate Transport Rate Coverage Balance Steps
Act 0.33 (0.19, 0.49) 0.67 (0.59, 0.76) 0.91 (0.86,0.95) 0.59 (0.52, 0.66) 24.92 (22.12,27.73)
ReAct 0.34 (0.20, 0.49) 0.72 (0.63,0.80) 0.92 (0.86, 0.97) 0.67 (0.61, 0.73) 24.08 (21.27, 26.89)
CoT 0.14 (0.06, 0.28) 0.59 (0.51, 0.67) 0.87 (0.81, 0.92) 0.62 (0.56,0.69) 28.4 (26.91, 29.97)
SmartLLM 0.11 (0.05, 0.23) 0.23 (0.13, 0.31) 0.91 (0.80, 0.96) 0.45 (0.37, 0.52) 29.87 (26.20, 30)
CoELA 0.25 (0.10, 0.36) 0.46 (0.35, 0.56) 0.76 (0.67, 0.85) 0.73 (0.67, 0.80) 28.93 (27.77,30)
LLaMAR 0.51 (0.36, 0.66) 0.85 (0.80, 0.91) 0.95 (0.91, 0.98) 0.83 (0.78, 0.86) 25.80 (23.72, 27.88)
(w/o vision)
LLaMAR 0.62 (0.46, 0.76) 0.87 (0.80, 0.93) 0.95 (0.91, 0.98) 0.82 (0.77, 0.87) 23.44 (20.88, 26.00)
(w/o exploration)
LLaMAR 0.66 (0.50, 0.78) 0.91 (0.81,0.96) 0.97 (0.93,0.99) 0.82 (0.75,0.87) 21.87 (18.76,24.23)
(w/ exploration)
Table 3: Comparison of evaluation metrics against baselines averaged across all tasks. More details about peculiar behaviors for the baselines can be found in Appendix E
@@
Roles of different modules in LLaMAR: To demonstrate the effectiveness of the various modules in our cognitive architecture, we performed ablation studies by evaluating performance metrics with each module removed individually. The results are summarized in Table 4. Using only the Actor module corresponds to the "Act" baseline, which demonstrates its fundamental capabilities in isolation but shows limited effectiveness without planning and correction due to relatively lower success and transport rates. Adding the Planner and Verifier modules improves performance, benefiting from better task planning and validation, increasing the overall SR and TR, and ensuring more effective task completion and even work distribution, as indicated by the increase in balance (B). However, in scenarios where the suggested action fails, the actor suggests the same action in the next decision step since it is not able to reason on why the action failed until the end of the planning horizon. Incorporating the
Modules Success Transport Coverage Balance Steps
Used Rate Rate
Actor 0.33 0.67 0.91 0.59 24.92
(0.19, 0.49) (0.59, 0.76) (0.86,0.95) (0.52, 0.66) (22.12,27.73)
Planner + 0.45 0.78 0.92 0.69 24.87
@@
LLaMAR 0.66 0.91 0.97 0.82 21.87
(0.50, 0.76) (0.81, 0.96) (0.93,0.99) (0.75, 0.87) (18.76, 26.43)
Corrector module with access to privileged information from an environment oracle significantly boosts performance, enhancing the SR, TR, and C, further improving B, and reducing L by approximately two time steps on average, consistent with the findings in (Arora and Kambhampati, 2023). This highlights the Corrector module’s importance in adjusting actions based on feedback, resulting in higher task success and more efficient task completion, albeit with reliance on oracle knowledge. Finally, the complete LLaMAR system, without privileged information, achieves SR, TR, C, and B values close to those of the oracle setup, with better L. This demonstrates the system’s robustness and effectiveness in a realistic setting. The Corrector module plays a crucial role in enabling agents to learn from past failures and avoid repeating actions, preventing task failures due to timeout. Despite lacking oracle knowledge, LLaMAR performs nearly as well as the oracle-enhanced setup. These results highlight the importance of each module in our cognitive architecture. Removing any module diminishes effectiveness, highlighting their essential roles in achieving state-of-the-art results.
@@
# agents Success Transport Coverage Balance Steps
Rate Rate
2 0.62 0.87 0.95 0.82 23.44
(0.46, 0.76) (0.80, 0.93) (0.91, 0.98) (0.77,0.87) (20.88, 26.00)
3 0.70 0.91 0.98 0.66 21.30
(0.55, 0.82) (0.85, 0.95) (0.95, 0.99) (0.61, 0.71) (18.60, 23.99)
4 0.68 0.90 0.99 0.62 22.83
(0.52, 0.79) (0.84, 0.94) (0.95, 0.99) (0.57, 0.68) (19.63, 25.69)
5 0.62 0.90 0.99 0.54 22.91
(0.46, 0.75) (0.85, 0.94) (0.97,1.00) (0.48, 0.59) (20.26, 25.57)
Table 5: LLaMAR with more agents
With 2 agents, we establish a solid baseline for success rate (SR) and transport rate (TR). Adding a third agent improves both SR and TR, indicating enhanced task completion and transportation efficiency. Coverage (C) also increases, suggesting better exploration and interaction with objects. However, balance (B), which measures the even distribution of tasks among agents, decreases with more agents. This drop highlights the challenge of ensuring equal contributions from all agents in a larger multi-agent system. While SR remains high, the balance metric drops significantly from 2 to 5 agents, indicating some agents do more work than others. The number of steps (L) taken to complete tasks generally decreases with more agents, showing improved efficiency but is less pronounced from 4 to 5 agents, suggesting diminishing returns on efficiency with each additional agent. In summary, adding more agents improves task performance and efficiency but introduces challenges in maintaining balanced contributions. Addressing this imbalance is crucial for refining multi-agent planning algorithms.

### v2

# LLaMAR: Long-Horizon Planning for Multi-Agent Robots in Partially Observable Environments

@@ Abstract
The ability of Language Models (LMs) to understand natural language makes them a powerful tool for parsing human instructions into task plans for autonomous robots. Unlike traditional planning methods that rely on domain-specific knowledge and handcrafted rules, LMs generalize from diverse data and adapt to various tasks with minimal tuning, acting as a compressed knowledge base. However, LMs in their standard form face challenges with long-horizon tasks, particularly in partially observable multi-agent settings. We propose an LM-based Long-Horizon Planner for Multi-Agent Robotics (LLaMAR), a cognitive architecture for planning that achieves state-of-the-art results in long-horizon tasks within partially observable environments. LLaMAR employs a plan-act-correct-verify framework, allowing self-correction from action execution feedback without relying on oracles or simulators. Additionally, we present MAP-THOR, a comprehensive test suite encompassing household tasks of varying complexity within the AI2-THOR environment. Experiments show that LLaMAR achieves a 30% higher success rate than other state-of-the-art LM-based multi-agent planners in MAP-THOR and Search & Rescue tasks. Code can be found at https://github.com/nsidn98/LLaMAR
**footnotetext: Equal Contribution.$\dagger$\$$\dagger$\$footnotetext: Now at Physical Intelligence.$\ddagger$\$$\ddagger$\$footnotetext: Work done outside Apple.
@@ 3 Background
Problem Setting: We consider a setting where multiple robots perform a series of tasks (a) such as cleaning a room or putting groceries in the fridge, in a home-like environment, and (b) rescuing missing personnel and putting out forest fires in a search & rescue environment (SAR). These tasks typically require long-horizon planning, involving around 100 low-level actions to reach the goal. Our objective is to compute plans for a team of robots to execute high-level language instructions, $I$. We formalize these tasks as partially observable Markov decision processes (POMDP) Puterman (1994); Kaelbling et al. (1998), denoted as $\langle N,\mathcal{I},\mathcal{S},\{\mathcal{O}_{i}\},\{\mathcal{A}_{i}\},\mathcal{P},\mathcal{G},T\rangle$. $N$ is the number of agents and $\mathcal{I}$ is the high-level language instruction set. Here, $s\in\mathcal{S}$ represents the joint state of all agents, and $o\in\mathcal{O}$ denotes the observation set for all agents. Particularly, $o_{i}\in\mathcal{O}_{i}$ is the observation set of agent $i$, that captures incomplete environment state information. $a\in\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\cdots\mathcal{A}_{N}$ represents the joint action space. The joint action space comprises of different categories of high-level actions $\mathcal{A}=\mathcal{A}_{NAV}\cup\mathcal{A}_{INT}\cup\mathcal{A}_{EXP}$, where $\mathcal{A}_{NAV}$ is the joint navigation action set, $\mathcal{A}_{INT}$ is the joint interaction actions which allow the agents to interact with objects, and $\mathcal{A}_{EXP}$ are the joint exploration actions which allow the agents to explore the environment. Examples of the high-level actions include PickUp(object) $\in\mathcal{A}_{INT}$ and NavigateTo(location) $\in\mathcal{A}_{NAV}$. Each high-level action is associated with a low-level primitive action (pre-trained RL, behavior cloned, or heuristic-based policy). These actions are executed synchronously by all agents at every high-level decision step. $\mathcal{P}(s^{\prime}|s,a)$ is the joint transition probability function that defines the probability of arriving at $s^{\prime}\in\mathcal{S}$ after taking joint action $a\in\mathcal{A}$ in $s\in\mathcal{S}$. $\mathcal{G}=\{g_{1},\cdots,g_{k}\}$ defines the subtasks that the agents need to perform to accomplish the language instruction task. $T$ is the length of the planning horizon.
@@
Environment: To simulate open-ended, long-horizon tasks that resemble everyday activities in a home-like environment, we use the AI2Thor simulator Kolve et al. (2017), which supports a diverse set of interactions and photorealistic rendering. Since our approach does not require any parametric training, it can potentially translate to other similar embodied environments like VirtualHome Puig et al. (2018), Habitat Savva et al. (2019); Szot et al. (2021); Puig et al. (2023), and ThreeDWorld Gan et al. (2021), possibly extending beyond household domains. Similarly, to simulate the search & rescue scenario, we create a custom Search & Rescue environment (SAR). More information about the MAP-THOR and SAR environments can be found in Appendix B and D respectively. We instantiate the problem with $N$ agents cooperating to accomplish a long-horizon rearrangement task Batra et al. (2020) in an indoor environment. The agents do not know the objects present in the environment a priori and are encouraged to explore the environment to gather more information to complete the task. Unlike previous solutions in similar settings Zhang et al. (2024); Kannan et al. (2023); Wang et al. (2024), we do not rely on an oracle/simulator to verify subtask completion. In prior works, a predefined conditional satisfiability check is used as subtask completion feedback.
@@ 4 Approach
Admissible Action parsing with Semantic Translation When LMs generate action plans, natural language outputs often fail to translate to executable high-level actions. This happens when the output does not match the predefined format or refers to unrecognized contextually similar objects. We use a cosine similarity method from (Huang et al., 2022), fine-tuning a pre-trained sentence-BERT (Reimers and Gurevych, 2019) to transform the free-form text into admissible high-level actions. Hyperparameters and additional details of the sentence transformer fine-tuning are provided in Appendix J.
@@
Motivation for Proposed Framework The specific order of the modules in LLaMAR is due to natural causal relationships in which environment feedback is received. We use the Planner as the first module because it allows LLaMAR to come up with an initial list of open subtasks that could be completed based on the current observation and past memory to satisfy the task. This list serves as a rough high-level plan. The actor then uses this information to suggest the necessary actions. The Corrector is used after the Actor module to identify reasons for failures in the execution of the actions suggested by the Actor . Note that the failure module is inert and only suggests corrective actions. Only the Actor module decides the final actions to be executed. This role distinction allows for clear reasoning on failures when they occur and lets the actor module focus on choosing actions. The Verifier is used after the action is executed to update the list of closed subtasks so that LLaMAR can be current with the progress toward the completion of the environment task. This allows the planner to update the list of open subtasks in the next step. In essence, the Planner and the Verifier ensure that the progress of the agents is tracked and the actor and the corrector ensure that the actions are executed successfully to advance towards completion of the task.
Multi-Agent features in LLaMAR While our method can be easily adapted to a single-agent setting, our design choice for the architecture was motivated to include the following multi-agent features:
  • Coordination through communication: Agents share their state information with the centralized LLaMAR modules to predict actions, enabling them to coordinate and avoid conflicts. This information sharing allows for the agents to cooperate and achieve the collective goal.
  • Dynamic Role Assignment: Agents are dynamically assigned roles based on the current task requirements and their capabilities. This flexibility allows LLaMAR to adapt to changing environments and task demands.
  • Hierarchical Task Decomposition: To handle the complexity of multi-agent planning, LLaMAR decomposes the action space by creating specific subgoals/subtasks available for any agent to assign itself (done by the actor module) based on the observation and current context. This decomposition reduces the overall search space and improves planning efficiency.
@@ 5 Experiments
MAP-THOR: To evaluate the performance of LLaMAR and benchmark other baseline methods, we create a benchmark dataset of tasks which we call MAP-THOR (Multi-Agent Planning tasks in AI2-THOR). While Smart-LLM Kannan et al. (2023) introduces a dataset of 36 tasks within AI2-Thor Kolve et al. (2017) classified by complexity, their tasks are limited to single floor plans. This limitation hinders testing the robustness of planners across different room layouts. Additionally, some tasks in their dataset cannot be performed by multiple agents, regardless of task division, such as Pick up the pillow, Open the laptop to turn it on, and Turn off the lamp.
@@
  • Implicit target location, item type, and quantity: Item types and their quantities along with the target location are implicitly defined. For example, Clear the floor by placing the items at their appropriate positions. The agent is expected to place items like pens, books, and laptops on the study table, and litter in the trash can.
Search & Rescue Environment (SAR): To showcase the effectiveness of LLaMAR with respect to explicit coordination in multi-agent settings, we evaluate LLaMAR in a partially observable search & rescue and fire relief environment in a grid world. Depending on the scene, there is a mix of missing people to be found, and wildfires to be stopped before they spread geographically. More details about the environment can be found in Appendix D.
  • Fire Extinguishing: Fires consist of expansive flammable regions with a fixed set of sources that propagate over time. The rate of fire spread is proportional to its intensity; higher intensities result in faster spread. Fires are categorized as either Class A or Class B, which are extinguished using water or sand, respectively. These extinguishing resources are sourced from reservoirs distributed across the environment.
  • Human Rescue: Each individual is initially located at an unknown position within the environment. The objective is to locate, rescue, and transport the individuals to a designated drop-off location, which is known beforehand. Transporting a person requires the coordinated effort of two agents simultaneously, who must carry them to the specified drop-off point.
@@
  • Transport Rate (TR): The fraction of subtasks completed within an episode, provides a finer granularity of task completion.
@@
  • Balance (B): The ratio between the minimum and maximum number of successful high-level actions executed by any agent that contributed towards task completion. We only check for a subset of high-level actions that must be executed to accomplish critical subtasks that lead to the successful completion of the language instruction task. If each agent $i$ out of $n$ agents completes $s_{i}$ successful tasks, the balance is defined as: $B:=\frac{\min\ \{s_{1},\cdots,s_{n}\}}{\max\{s_{1},\cdots,s_{n}\}+\epsilon}$. This measures how evenly the work is distributed among agents. A balance of zero indicates at least one agent performed no successful high-level actions, while a balance of one indicates all agents performed the same number of successful high-level actions. Here $\epsilon=1e-4$ is a small number to avoid division by zero.
  • Average steps (L): The number of high-level actions taken by the team to complete the task, capped at $L=30$ in our experiments. If the task is not completed within $L$ steps, the episode is deemed a failure. Note that the metric $L$ is presented in the table providing the complete results, located in Appendix F.
For all the metrics, we report the means along with the $95\%$ confidence interval across all the tasks (refer Appendix F for complete results). Since SR is a binomial metric, we report the Clopper-Pearson Interval as the confidence interval.
Baselines
For a fair comparison with our method, we make modifications to the baselines to make them work in partially observable settings with limited reliance on the simulator. More details about implementations can be found in Appendix H.
@@
  • ReAct Yao et al. (2023): We use a ReAct-style prompting to let the LMs reason after suggesting high-level actions and possibly suggest ways to correct any failures.
  • SmartLLM Kannan et al. (2023): We modify the official codebase to only include information from the local observations of the agents instead of assuming full observability.
  • CoELA Zhang et al. (2024): We modify the list of available high-level actions to include all possible valid combinations of actions with interactable objects in the agent’s local observation. As the scene becomes more cluttered, this list and the prompt becomes combinatorially longer. In the original implementation, the list of available actions is filtered based on the feasibility of the actions as suggested by a conditional checker.
It should be noted that Act, Chain-of-Thought, ReAct, and SmartLLM are all CMAS frameworks where CoELA follows the DMAS framework.
@@ 6 Results and Discussion
Choice of the underlying LM: To understand the impact of the underlying LM’s quality on decision-making, we initially experimented with different LMs on MAP-THOR. Specifically, we utilize both the language-only and vision-language models of GPT-4 OpenAI et al. (2023), IDEFICS-2 Laurençon et al. (2023, 2024), LLaVA Liu et al. (2023b, a), and CoGVLM Hong et al. (2023a). Among these, GPT-4, when used solely with text inputs, exhibits the poorest performance. This is attributed to the agents’ inability to reason about visual observations, which is particularly detrimental for the Corrector module. Substituting GPT-4V with other vision-language models results in a decline in performance (refer Table 3) and hence we use GPT-4V as the underlying VLM while comparing to the baselines.
@@
Baseline Comparisons: Table 3 compares our method, LLaMAR, with other baselines in a 2-agent scenario using GPT-4V as the underlying VLM. Act and ReAct show similar performance, with Act struggling due to its lack of strategic planning or correction, and ReAct performing slightly better by dynamically adjusting actions based on reasoning on immediate feedback. CoT’s performance declines with longer planning horizons due to its inability to maintain coherence over extended planning sequences, consistent with findings in (Stechly et al., 2024), showing its effectiveness only with highly specific prompts. SmartLLM, operating in a plan-and-execute paradigm, generates impractical plans with issues like infinite loops and failure to handle low-level action failures, leading to lower success rates and poor transport metrics. It also tends to hallucinate objects. CoELA, using a decentralized multi-agent system (DMAS), performs poorly due to large input prompts and struggles to select the correct action from numerous choices. Its decentralized decision-making is less efficient than the centralized multi-agent system (CMAS) used by LLaMAR. Previous research (Chen et al., 2023) confirms CMAS frameworks are more effective than DMAS frameworks. Overall, our method, LLaMAR, benefits from its modular cognitive architecture, which integrates planning, acting, correcting, and verifying through distinct LLM roles, resulting in superior performance across various evaluation metrics. By avoiding reliance on privileged information and incorporating a robust exploration strategy that allows it to scout for objects that are not initially visible, LLaMAR ensures higher success rates and balanced task execution among agents.
@@
Roles of different modules in LLaMAR: To demonstrate the effectiveness of the various modules in our cognitive architecture, we performed ablation studies by evaluating performance metrics with each module removed individually. The results are summarized in Table 3. Using only the Actor module corresponds to the "Act" baseline, which demonstrates its fundamental capabilities in isolation but shows limited effectiveness without planning and correction due to relatively lower success and transport rates. Adding the Planner and Verifier modules improves performance, benefiting from better task planning and validation, increasing the overall SR and TR, and ensuring more effective task completion and even work distribution, as indicated by the increase in balance (B). However, in scenarios where the suggested action fails, the actor suggests the same action in the next decision step since it is not able to reason why the action failed until the end of the planning horizon. Incorporating the Corrector module, with access to privileged information from an environment oracle, significantly boosts performance, enhancing the SR, TR, C, and further improving B, consistent with the findings in (Arora and Kambhampati, 2023). This highlights the Corrector module’s importance in adjusting actions based on controller feedback, resulting in higher task success and more efficient task completion, albeit with reliance on oracle knowledge. Finally, the complete LLaMAR system, without privileged information, achieves SR, TR, C, and B values close to those of the oracle setup. This demonstrates the system’s robustness and effectiveness in a realistic setting. The Corrector module plays a crucial role in enabling agents to learn from past failures and avoid repeating actions, preventing task failures due to timeout. Despite lacking oracle knowledge, LLaMAR performs nearly as well as the oracle-enhanced setup. These results highlight the importance of each module in our cognitive architecture. Removing any module diminishes effectiveness, highlighting their essential roles in achieving state-of-the-art results.
[…]
      Act     GPT-4V     0.33 0.67 0.91 0.59
    ReAct     GPT-4V     0.34 0.72 0.92 0.67
    CoT     GPT-4V     0.14 0.59 0.87 0.62
    SmartLLM     GPT-4V     0.11 0.23 0.91 0.45
    CoELA     GPT-4V     0.25 0.46 0.76 0.73
      LLaMAR     GPT-4     0.51 0.85 0.95 0.83
    LLaMAR     LLaVA     0.54 0.84 0.91 0.75
    LLaMAR     IDEFICS-2     0.57 0.86 0.94 0.78
    LLaMAR     CogVLM     0.61 0.89 0.95 0.80
    LLaMAR     GPT-4V     0.62 0.87 0.95 0.82
    (w/o expl)
    LLaMAR     GPT-4V     0.66 0.91 0.97 0.82
    (w/ expl)
Table 2: Comparison of evaluation metrics against baselines averaged across all tasks for the 2-agent MAP-THOR scenarios. The complete table with confidence intervals can be found in Appendix F. More details about peculiar behaviors for the baselines can be found in Appendix H.
      Modules     SR $\uparrow$ TR $\uparrow$ C $\uparrow$ B $\uparrow$
    Used
      Actor     0.33 0.67 0.91 0.59
    Planner +     0.45 0.78 0.92 0.69
    Actor +
@@
    Planner +     0.66 0.91 0.97 0.82
    Actor +
    Corrector +
    Verifier +
Table 3: Performance in the 2-agent scenarios in MAP-THOR obtained by ablating different modules in LLaMAR with GPT-4V as the underlying LM
@@
      # of     MAP-THOR     SAR
[…]
      1     0.37     0.67 0.87 1.00     0.28 0.75 0.86 1.00
    2     0.62     0.87 0.95 0.82     0.44 0.86 0.94 0.91
    3     0.70     0.91 0.98 0.66     0.68 0.92 0.96 0.80
    4     0.68     0.90 0.99 0.62     0.72 0.94 0.98 0.78
    5     0.62     0.90 0.99 0.54     0.74 0.96 1.00 0.73
Table 4: LLaMAR with various number of agents in the scenario in both MAP-THOR and SAR environments
Increasing the number of agents in the environment shows distinct trends in our method’s performance metrics for both MAP-THOR and SAR environments (refer Table 4). In MAP-THOR, with two agents, we establish a solid baseline for success rate (SR) and transport rate (TR), which is similarly reflected in the SAR environment. Adding a third agent improves both SR and TR in both environments, indicating enhanced task completion and transportation efficiency. Coverage (C) also increases, suggesting better exploration and interaction with objects across both environments. There is a slight decrease in SR and TR when the number of agents increases from 3 to 5 in the MAP-THOR environment. The decrease in these metrics can be attributed to the rooms in MAP-THOR becoming crowded with 4 and 5 agents hence blocking the agents from navigating without colliding with other agents. But this phenomenon is not seen in the SAR environment which is comparatively more spacious and navigable. However, balance (B), which measures the even distribution of tasks among agents, decreases with more agents. This drop highlights the challenge of ensuring equal contributions from all agents in a larger multi-agent system. While SR remains high, the balance metric drops significantly from 2 to 5 agents, indicating some agents do more work than others. In summary, adding more agents improves task performance and efficiency but introduces challenges in maintaining balanced contributions. Addressing this imbalance is crucial for refining multi-agent planning algorithms.